• request by Daniel Valuch for CV assistance to purge the cooling water system for the SRF amplifiers (in B199). To be checked with Xavier Decrand (retour 26 Jan).
  • HIE tunnel visit by Rustan Calming EN-EL: future cable work.
  • HIE tunnel BCAM system (MATHILDE) disconnected (Antje Behrens, Alexandre Beynel: BE-GM Survey). Survey planning OK for CM1 dismount (start Mon 2 March).
  • Test for fitting of the lifting pins for the CM handling tool today in the HIE tunnel (Mathieu Therasse).
  • Replacement of the RC4.FC90 unit foreseen today has been postoned RC4 was under vacuum. VSC (Guillermo Fernandez) wants to be present during the intervention. Vacuum acceptance test to be performed? Opened the N2 bottles for venting process (Erwin). BI wondering who to contact for OP related issues.